# pnasresearcharticle Xingying Liu

### v1

@@ Abstract
Cluster analysis is a fundamental research issue in statistics and machine learning. In many modern clustering methods, we need to determine whether two subsets of samples come from the same cluster. Since these subsets are usually generated by certain clustering procedures, the deployment of classic two-sample tests in this context would yield extremely smaller -values, leading to inflated Type-I error rate. To overcome this bias, we formally introduce the two-cluster test issue and argue that it is a totally different significance testing issue from conventional two-sample test. Meanwhile, we present a new method based on the boundary points between two subsets to derive an analytical -value for the purpose of significance quantification. Experiments on both synthetic and real data sets show that the proposed test is able to significantly reduce the Type-I error rate, in comparison with several classic two-sample testing methods. More importantly, the practical usage of such two-cluster test is further verified through its applications in tree-based interpretable clustering and significance-based hierarchical clustering.
@@
Clustering analysis Two-sample test Hypothesis testing Statistical significanceIn this paper, we formally introduce a new hypothesis testing problem in the context of cluster analysis, namely two-cluster test. Given two disjoint subsets of samples from a data set that is composed of one or more clusters, the objective of such a test is determine whether they belong to the same cluster or not.
@@
One may argue that we can apply existing two-sample test methods (7) to solve the above issue. However, two-sample tests control the Type-I error rate with respect to the mean difference when two subsets are defined a priori (8, 9). In above applications, the subsets of samples are instead generated by clustering methods. Therefore, the direct application of a classical two-sample test is not feasible, which can yield an extremely inflated Type-I error rate, as illustrated in Fig.Β 1 and observed in (6, 4). In order to apply the two-sample test to obtain valid inference result in this context, one may seek to employ the selective inference technique (8, 9). However, the selective inference method is highly dependent on the clustering algorithm employed (e.g. -means (8), hierarchical clustering (9)), making it difficult to become a general significance test.
In significance-based hierarchical clustering (4, 5, 6), such two-cluster test issue is properly tackled via simulation-based approach. That is, Monte Carlo sampling is used to generate simulated data sets whose cluster assignments are compared to those of the original data to compute empirical -values. Although intuitive, this procedure incurs substantial computational cost and it cannot provide an analytical -value.
Ideally, we would like to be able to yield an analytical -value for any two subsets of samples under the assumption that they belong to the same cluster. However, there are still no research efforts towards this direction and such two-cluster test issue is rarely known in the literature.
To address this gap, we formally formulate the two-cluster test issue and present a new significance testing procedure. The new significance testing method is called BTCT (Boundary-based Two-Cluster Test). Specifically, BTCT defines boundary points as samples that are mutual -nearest neighbors between the two subsets. Under the null hypothesis that the two subsets originate from the same cluster, for each boundary point, the number of its -nearest neighbors that fall into the same subset is used as the test statistic, which follows a Binomial(, 0.5) distribution. All boundary point -values are then combined via meta-analysis to obtain a census -value.
@@
In Fig.Β 2(a), we first identify boundary points between the two subsets, and then find their nearest neighbors, as shown in Fig.Β 2(a2). For each boundary point, we count how many neighbors belong to the same subset to calculate the test statistic. Under the null hypothesis that subsets and are from the same cluster, the statistic follows a distribution. As shown in Fig.Β 2(a3), all the nearest neighbors of each boundary point come from the same subset, yielding small -values of 0.0078. The combined -value is by using Fisherβs method, which is less than the significance threshold , leading us to reject the null hypothesis and indicating that two subsets do not belong to the same cluster.
@@
To benchmark our method, we compare it against three classic two-sample tests, all implemented using the Torch Two Sample package 111https://github.com/josipd/torch-two-sample: FR (20), MMD (21) and Energy (22). For these methods, we compute the -values using permutation testing with 1000 permutations.
The FR test constructs an minimum spanning tree (MST) over the pooled samples and counts the number of edges that connect samples from different groups, thereby testing whether the two sets of samples originate from distinct distributions.
@@
For each data set, we apply min-max normalization to ensure consistency in scale, then split each individual cluster into two equally sized subsets based on the median of first feature values and test whether these two subsets are still identified as belonging to the same cluster. For the MMD test, we use a Gaussian kernel with its bandwidth parameter being equal to the median of pairwise euclidean distances. In BTCT, we set the number of neighbors for each boundary point to . The accuracy is calculated as the ratio of subset pairs that are correctly identified based on their -values. Specifically, a subset pair is correctly identified if the corresponding -value is less than (or larger) the significance threshold of 0.05 and they come from the different clusters (or the same cluster). Detailed results are presented in Fig.Β 3.
@@
Currently, tree-based interpretable clustering methods (5, 6, 4) typically assess the statistical significance of candidate splits using two-sample tests. However, such tests often yield excessively small -values, leading to potentially unreliable splits. Therefore, we apply BTCT instead of two-sample test in this context to identify truly meaningful splits to guide the tree growth, thereby avoiding the possibility of over- or under-clustering. In the following, we illustrate this application with an example.
As depicted in Fig.Β 5, for each feature, samples are sorted in a non-decreasing order, and all possible split points that can partition the data into two subsets are examined. For each candidate split, we examine whether two corresponding subsets belong to the same cluster or not based on either two-cluster test or two-sample test. The smallest -value across all candidate splits is then compared to the significance threshold : if it falls below this threshold, the corresponding split is considered to be significant to form a branch node; otherwise, the current node is treated as a leaf node. Recursion prioritizes the left child of any branch node, which serves as the new input for this tree growth process, continuing until no leaf nodes can be further divided.
@@
We assess the explainability of clustering results by comparing three metrics: the average depth of the tree (avgDepth), the maximal depth of the tree (maxDepth), and number of leaf nodes (nLeaf). Simpler tree structures with fewer levels and leaf nodes are generally easier to interpret, which indicates better explainability.
@@
One critical issue in divisive hierarchical clustering is when should we terminate the split operation. If it is terminated earlier or later, then under or over-clustering will occur. Hence, people begin to employ hypothesis testing methods to assess the statistical significance of each division, thereby yielding a right cluster number estimation (1, 2, 3). The idea is similar to that used in tree-based interpretable clustering, we can assess whether two divided subsets belong to the same cluster via two-cluster test or two-sample test. An example is given in Fig.Β 8, where we use our method to determine whether the corresponding division is statistically significant.
@@
Fig.Β 9 compares the predicted number of clusters by four methods across 12 data sets. The red shaded region is the interval for each data set, where is the number of ground-truth clusters. The predicted cluster number by BTCT is close to this region on most data sets, slight overestimation occurs due to the impurity of the subsets created by hierarchical clustering method based on Ward linkage at each split. In contrast, the baseline methods often yield extremely small -values for candidate divisions, leading to over-clustering and overestimation of cluster number.
@@ Discussion
Here we introduce the two-cluster test, a significance testing issue that as been overlooked in the field of cluster analysis. We highlight the importance of such a new test and demonstrate that classic two-sample tests fail to control the Type-I error rate when the objective is to determine whether two subsets of samples come from the same cluster. More importantly, the two-cluster test is the key issue in modern clustering algorithms such as tree-based interpretable clustering algorithms and significance-based hierarchical clustering algorithms.
There are still many interesting issues that remain unaddressed. Firstly, the performance of proposed BTCT method is still far from being satisfactory. Hence, more effective methods should be further developed. Secondly, new clustering algorithms that are equipped with two-cluster test should be developed so as to demonstrate its practical usage. Finally, more application scenarios for the two-cluster test should be investigated, such as hierarchical community detection (23).

### v2

@@ Abstract
Cluster analysis is a fundamental research issue in statistics and machine learning. In many modern clustering methods, we need to determine whether two subsets of samples come from the same cluster. Since these subsets are usually generated by certain clustering procedures, the deployment of classic two-sample tests in this context would yield extremely smaller -values, leading to inflated Type-I error rate. To overcome this bias, we formally introduce the two-cluster test issue and argue that it is a totally different significance testing issue from conventional two-sample test. Meanwhile, we present a new method based on the boundary points between two subsets to derive an analytical -value for the purpose of significance quantification. Experiments on both synthetic and real data sets show that the proposed test is able to significantly reduce the Type-I error rate, in comparison with several classic two-sample testing methods. More importantly, the practical usage of such a two-cluster test is further verified through its applications in tree-based interpretable clustering and significance-based hierarchical clustering.
@@
Clustering analysis Two-sample test Hypothesis testing Statistical significanceIn this paper, we formally introduce a new hypothesis testing problem in the context of cluster analysis, namely two-cluster test. Given two disjoint subsets of samples from a data set that is composed of one or more clusters, the objective of such a test is to determine whether they belong to the same cluster or not.
@@
One may argue that we can apply existing two-sample test methods (7) to solve the above issue. However, two-sample tests control the Type-I error rate with respect to the mean difference when two subsets are defined a priori (8, 9). In the above applications, the subsets of samples are instead generated by clustering methods. Therefore, directly applying a classical two-sample test is not feasible, which can yield an extremely inflated Type-I error rate, as illustrated in Fig.Β 1 and observed in (6, 4). In order to apply the two-sample test to obtain a valid inference result in this context, one may seek to employ the selective inference technique (8, 9). However, the selective inference method is highly dependent on the clustering algorithm employed (e.g. -means (8) and hierarchical clustering (9)), making it difficult to become a general significance test.
In significance-based hierarchical clustering (1, 2, 3), such two-cluster test issue is properly tackled via a simulation-based approach. That is, Monte Carlo sampling is used to generate simulated data sets whose cluster assignments are compared to those of the original data to compute empirical -values. Although intuitive, this procedure incurs substantial computational cost and it cannot provide an analytical -value.
Ideally, we would like enable an analytical -value for any two subsets of samples under the assumption that they belong to the same cluster. However, there are still no research efforts towards this direction and such a two-cluster test issue is rarely known in the literature.
To fill this gap, we formally formulate the two-cluster test issue and present a new significance testing procedure. The new significance testing method is called BTCT (Boundary-based Two-Cluster Test). Specifically, BTCT defines boundary points as samples that are mutual -nearest neighbors between the two subsets. Under the null hypothesis that the two subsets originate from the same cluster, for each boundary point, the number of its -nearest neighbors that fall into the same subset is used as the test statistic, which follows a Binomial(, 0.5) distribution. All boundary point -values are then combined via meta-analysis to obtain a census -value.
@@
In Fig.Β 2(a), we first identify boundary points between the two subsets, and then find their nearest neighbors, as shown in Fig.Β 2(a2). For each boundary point, we count how many neighbors belong to the same subset to calculate the test statistic. Under the null hypothesis that the subsets and are from the same cluster, the statistic follows a distribution. As shown in Fig.Β 2(a3), all the nearest neighbors of each boundary point come from the same subset, yielding small -values of 0.0078. The combined -value is by using Fisherβs method, which is less than the significance threshold , leading us to reject the null hypothesis and indicating that the two subsets do not belong to the same cluster.
@@
To validate our method, we compare it against three classic two-sample tests, all implemented using the Torch Two Sample package 111https://github.com/josipd/torch-two-sample: FR (20), MMD (21) and Energy (22). For these methods, we compute the -values using permutation testing with 1000 permutations.
The FR test constructs a minimum spanning tree (MST) over the pooled samples and counts the number of edges that connect samples from different groups, thereby testing whether the two sets of samples originate from distinct distributions.
@@
For each data set, we apply min-max normalization to ensure consistency in scale, then split each individual cluster into two equally sized subsets based on the median of the first feature values and test whether these two subsets are still identified as belonging to the same cluster. For the MMD test, we use a Gaussian kernel with its bandwidth parameter being equal to the median of pairwise euclidean distances. In BTCT, we set the number of neighbors for each boundary point to . The accuracy is calculated as the ratio of subset pairs that are correctly identified based on their -values. Specifically, a subset pair is correctly identified if the corresponding -value is less than (or larger than) the significance threshold of 0.05 and they come from the different clusters (or the same cluster). Detailed results are presented in Fig.Β 3.
@@
Currently, tree-based interpretable clustering methods (5, 6, 4) typically assess the statistical significance of candidate splits using two-sample tests. However, such tests often yield excessively small -values, leading to potentially unreliable splits. Therefore, we apply BTCT instead of a two-sample test in this context to identify truly meaningful splits to guide the tree growth, thereby avoiding the possibility of over- or under-clustering. In the following, we illustrate this application with an example.
As depicted in Fig.Β 5, for each feature, samples are sorted in a non-decreasing order, and all possible split points that can partition the data into two subsets are examined. For each candidate split, we examine whether the two corresponding subsets belong to the same cluster or not based on either a two-cluster test or a two-sample test. The smallest -value across all candidate splits is then compared to the significance threshold : if it falls below this threshold, the corresponding split is considered to be significant to form a branch node; otherwise, the current node is treated as a leaf node. Recursion prioritizes the left child of any branch node, which serves as the new input for this tree growth process, continuing until no leaf nodes can be further divided.
@@
We assess the explainability of clustering results by comparing three metrics: the average depth of the tree (avgDepth), the maximal depth of the tree (maxDepth), and the number of leaf nodes (nLeaf). Simpler tree structures with fewer levels and leaf nodes are generally easier to interpret, which indicates better explainability.
@@
One critical issue in divisive hierarchical clustering is when should we terminate the split operation. If it is terminated earlier or later, then under- or over-clustering will occur. Hence, people begin to employ hypothesis testing methods to assess the statistical significance of each division, thereby yielding a right cluster number estimation (1, 2, 3). The idea is similar to that used in tree-based interpretable clustering, we can assess whether two divided subsets belong to the same cluster via a two-cluster test or a two-sample test. An example is given in Fig.Β 8, where we use our method to determine whether the corresponding division is statistically significant.
@@
Fig.Β 9 compares the predicted number of clusters by four methods across 12 data sets. The red shaded region is the interval for each data set, where is the number of ground-truth clusters. The predicted cluster number by BTCT is close to this region on most data sets, and slight overestimation occurs due to the impurity of the subsets created by the hierarchical clustering method based on Ward linkage at each split. In contrast, the baseline methods often yield extremely small -values for candidate divisions, leading to over-clustering and overestimation of cluster number.
@@ Discussion
We introduce the two-cluster test, a significance testing issue that has been overlooked in the field of cluster analysis. We highlight the importance of such a new test and demonstrate that classic two-sample tests fail to control the Type-I error rate when the objective is to determine whether two subsets of samples come from the same cluster. More importantly, the two-cluster test is the key issue in modern clustering algorithms such as tree-based interpretable clustering algorithms and significance-based hierarchical clustering algorithms.
There are still many interesting issues that remain unaddressed. Firstly, the performance of the proposed BTCT method is still far from being satisfactory. Hence, more effective methods should be further developed. Secondly, new clustering algorithms that are equipped with two-cluster test should be developed so as to demonstrate its practical usage. Finally, more application scenarios for the two-cluster test should be investigated, such as hierarchical community detection (23).